# Goal-Oriented Multiple Access Connectivity for Networked Intelligent Systems

### v1

@@ IV-D Decision Making Criterion
To obtain the threshold criterion $v_{{\rm th},k}^{(n)}$, $\forall k$, $\forall n\in\mathcal{A}_{k}$, for the self-decision scheme proposed in II-A, the $k$-th ISA needs to timely compute $\alpha_{k,n}^{*}$, $\forall k,n$, according to Section IV-B. To this end, the ISA measures its distance from all NMAs thanks to the received query signals and has knowledge of $\mathcal{K}_{n}$, the other ISAs’ location distribution, content usefulness distributions, and observation accuracy. We derive the following upper bound for $\gamma_{k,n}$ in (II-B) by applying the Markov’s inequality:
$\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}\leq$
$~{}~{}~{}\sum_{m=1}^{M}\underset{\mathcal{C}_{k}}{\mathbb{E}}\!\left\{\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\!\left(\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!\!\left(\Lambda_{k^{\prime},m}^{(n)}\right)^{\!3}}\,\Omega_{k^{\prime\prime},m}^{(n)}\right\}\!.$ (19)
After equalizing $\gamma_{k,n}^{(\mathcal{C},\mathcal{I})}$ to its upper bound and using Algorithm 1, $\alpha_{k,n}^{*}$ is derived for the $k$-th ISA and $n$-th attribute at any arbitrary service period. Eventually, we arrive
@@
where $F_{V}(v_{k,j}^{(n)})$ indicates the CDF of the content meta-value. Also, $\beta_{k,n}$ shows the scheme-dependent probability that the $k$-th ISA generates contents for the $n$-th attribute, given by
@@ V Simulation Results
We investigate the performance of the proposed self-decision scheme in the system, satisfying the goal with $U_{\rm GoE}=1$. By default, we set $K=2$, $M=1$, $I_{n}=5$, $\forall n$, $p=0.2$ (fast-changing regime), and $\chi_{k,n}=0.8$, $\forall k,n$. ISAs are distributed uniformly in an area of $50\times 50\,[\text{m}^{2}]$, and NMAs are located in a formation to cover all ISAs. The assignment of a meta-value to each content follows a Gaussian distribution with mean $0.5$ and variance $0.2$. Besides, $P_{X}(\cdot)$ denotes the same distribution, but with mean 00 and variance $5$. Moreover, $g_{1}$, $g_{2}$, and $g_{3}$ have exponential forms, while $h$ and $f_{\rho}$ are linear. Also, the path-loss exponent $a=3.8$, $\sigma_{m}^{2}=-40\,[\text{dBm}]$, and $\gamma_{\rm th}=25\,[\text{dB}]$.
Fig. 2 illustrates the system’s effectiveness performance using the obtained optimal activation probabilities and weight factors of the considered features. The semantics-aware scheme empowered with the self-decision scheme represents the highest GoE for communicated contents from different viewpoints. This performance comes with a smaller discrepancy error, higher useful content acquisition, and fewer drop-offs compared to the other schemes. For instance, in Fig. 2 (a), the semantics-aware scheme offers on average $2\%$ and $2.19\%$ higher GoE with $25.13\%$ and $6.09\%$ lower drop-offs of contents than those of the uniform and change-aware schemes, respectively. Besides, the self-decision scheme adds $21.38\%$, $20.6\%$, and $27.33\%$ to the GoE and decreases the transmission rate by $42.2\%$, $67.21\%$, and $48.3\%$ for the semantics-aware, uniform, and change-aware schemes, respectively, compared to the case with no decision-making, where $\alpha_{k,n}=\beta_{k,n}$, $\forall k,n$. In a nutshell, the joint self-decision scheme and the semantics-aware scheme provide on average up to $29.52\%$ higher GoE, $25.13\%$ fewer drop-offs, and $67.21\%$ lower transmission rate.
[…]
[scale=0.215]Results/Dimension_NEW.eps
[…]
[scale=0.215]Results/NMA_ISASize_NEW.eps
[…]
[scale=0.215]Results/AttributeSize_NEW.eps
[…]
[scale=0.215]Results/StateSize_NEW.eps
Figure 2: The interplay between GoE and (a) the coverage area dimension, (b) the number of ISAs and NMAs, (c) the number of queried attributes, and (d) the size of attribute state space.
Also, as in Fig. 2 (a), widening the coverage area reduces the GoE under a fixed number of agents since the transmission success probability dwindles with distance. Moreover, Fig. 2 (b) shows that increasing $K$ and/or $M$ enhances the GoE. Although a larger $K$ boosts the coverage density, hence the success probability, it gradually raises the number of interfering agents. Therefore, the offered GoE is saturated for large enough $K$. According to Fig. 2 (c), the GoE decreases if the NMAs require more attribute levels, since discrepancy errors are accumulated at each service period. In Fig. 2 (d), increasing $I_{n}$, $\forall n$, brings in higher EDE and lower ERC with respect to the EUoC, while counterbalancing each other’s impacts for larger $I_{n}$. For the change-aware scheme with $I_{n}<6$, the ERC is larger than the other features, as the system works in a fast-changing regime.
@@ VI Conclusion
We developed a self-decision goal-oriented multiple access scheme for networked intelligent systems. Aiming to maximize the GoE, we obtained the optimal activation probabilities and threshold criteria for decision-making under different content acquisition schemes. Our results showed that our proposal could provide on average up to $29.52\%$ higher GoE, $25.13\%$ fewer drop-offs, and $67.21\%$ lower transmission rate.

### v2

@@ IV-D Decision Making Criterion
To obtain the threshold criterion $v_{{\rm th},k}^{(n)}$, $\forall k$, $\forall n\in\mathcal{A}_{k}$, for the self-decision scheme proposed in II-A, the $k$-th ISA needs to timely compute $\alpha_{k,n}^{*}$, $\forall k,n$, according to Section IV-B. To this end, the ISA measures its distance from all NMAs thanks to the received query signals and has knowledge of $\mathcal{K}_{n}$, the other ISAs’ location distribution, update usefulness distributions, and observation accuracy. We derive the following upper bound for $\gamma_{k,m,n}$ in (II-B) by applying the Markov’s inequality
$\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}\!\leq\!\underset{\mathcal{C}_{k}}{\mathbb{E}}\!\left\{\sum_{k^{\prime}\in\mathcal{C}_{k}}\sum_{k^{\prime\prime}\in\mathcal{I}}\!\dfrac{\prod_{k^{\prime}_{0}\in\mathcal{C}_{k}}\Lambda_{k^{\prime}_{0},m}^{(n)}\prod_{k^{\prime\prime}_{0}\in\mathcal{I}}\Omega_{k^{\prime\prime}_{0},m}^{(n)}}{\!\left(\Lambda_{k^{\prime},m}^{(n)}\!+\!\Omega_{k^{\prime\prime},m}^{(n)}\right)\!\!\left(\Lambda_{k^{\prime},m}^{(n)}\right)^{\!3}}\,\Omega_{k^{\prime\prime},m}^{(n)}\!\right\}\!.$ (18)
After equalizing $\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}$ to its upper bound and using Algorithm 1, $\alpha_{k,n}^{*}$ is derived for the $k$-th ISA and $n$-th attribute at any arbitrary service interval. Eventually, we arrive
@@
where $F_{V}(v_{k,j}^{(n)})$ indicates the CDF of the update’s meta value. Also, $\beta_{k,n}$ shows the scheme-dependent probability that the $k$-th ISA generates updates for the $n$-th attribute, given by
@@ V Simulation Results
We investigate the performance of the proposed self-decision scheme for satisfying the goal with ${\rm{EUU}}_{\rm min}=0.1$. We consider $K=10$, $M=4$, $I_{n}=10$, $p^{\prime}_{n}=0.2$, $\forall n$, $|\widehat{\mathcal{A}}_{t}|=10$, $M_{t}=\lceil\frac{M+1}{2}\rceil=3$, $\forall t$, and $q_{k,n}=0.8$, $\forall k,n$. We also randomly select a subset of the ISAs constructing $\mathcal{K}_{n}$, $\forall n$. Additionally, we assume that the distance between each pair of ISA and NMA having a $7\,[\text{m}]$ height follows a standard normal distribution with standard deviation $60\,[\text{m}]$, the path-loss exponent $a=3.8$, $\sigma_{m}^{2}=-120\,[\text{dBm}]$, and $\gamma_{\rm th}=10\,[\text{dB}]$. The assignment of a meta value to each update is based on a beta distribution with both shape parameters equal to $2$. Moreover, $g_{1}$, $g_{2}$, and $g_{3}$ have exponential forms, while $h$ and $f_{\rho}$ are linear and cubic functions, respectively. To simplify, we assume $\alpha_{k,n}=\alpha_{k^{\prime},n^{\prime}}$ and accordingly $v_{{\rm th},k}^{(n)}=v_{{\rm th},k^{\prime}}^{(n^{\prime})}$, $\forall n,n^{\prime},k,k^{\prime}$.
Fig. 2 illustrates the effect of the meta value threshold, i.e., $v_{{\rm th},k}^{(n)}$, $\forall n,k$, hence activation probability, on the effectiveness. The optimization objective of the effectiveness problem, i.e., $\mathcal{G}$ in $\mathcal{P}_{2}$, and its EUU constraint are plotted versus $v_{{\rm th},k}^{(n)}$ based on different update acquisition schemes and $\mathbf{W}=[0.5,0.5]^{T}$ using Monte Carlo simulations with $500$ iterations. Despite almost identical $\mathcal{G}$, the semantics-aware scheme offers the highest EUU on average. To satisfy ${\rm{EUU}}_{\rm min}$, as shown in Fig. 2, the semantics-aware, uniform, and change-aware schemes require transmission rates of $\alpha_{k,n}=0.6$, $0.61$, and $0.62$, $\forall k,n$, respectively, under the first root, while they generate updates with rates of $0.91$, $1$, and $0.8$. Thus, we infer that the semantics-aware scheme poses, averagely, $6.4\%$ less channel load than the uniform scheme and $5\%$ more than the change-aware one.
To assess the self-decision scheme, we apply Algorithm 1 to derive the set of optimal activation probabilities based on different schemes and the assumed $\mathbf{W}$. In this case, we reach $\alpha_{k,n}=0.64$, $0.64$, and $0.67$, $\forall n,k$, for the semantics-aware, uniform, and change-aware schemes, respectively. Compared to the derived values from Fig. 2, the accuracy of the self-decision scheme is more than $92\%$. More importantly, ${\rm{EUU}}_{\rm min}$ is not satisfied without our self-decision scheme.
Figure 2: The objective of $\mathcal{P}_{2}$ and its constraint versus the meta value threshold.
The interplay between $\mathcal{G}$ and the number of states in the DTMC, i.e., $I_{n}$, $\forall n$, is depicted in Fig. 3 based on different values of $|\widehat{\mathcal{A}}_{t}|$, $\forall t$. Therein, we plot the lowest value of the objective that meets ${\rm{EUU}}_{\rm min}$. We see that increasing $I_{n}$ results in higher $\mathcal{G}$ due to the increase of $P_{{\rm e},n}$ and EDE, accordingly, concerning (4). Additionally, higher values of $|\widehat{\mathcal{A}}_{t}|$ boost ERC according to (7), leading to higher $\mathcal{G}$. Thus, the fewer attributes queried within a service interval, the higher the effectiveness.
Figure 3: The interplay between the objective of $\mathcal{P}_{2}$ and the number of states in the DTMC for different sizes of required attributes.
@@ VI Conclusion
We developed a self-decision goal-oriented multiple access scheme for networked intelligent systems. Aiming to maximize the GoE, we obtained the optimal activation probabilities and threshold criteria for decision-making under different update acquisition schemes. Analyzing how different parameters affect the effectiveness, our simulation results showed that our proposal could achieve at least $92\%$ of the optimal performance.